What is the frequency of a photon with a wavelength of 1040 nm. Report your answers to three significant digits.

The frequency is = _________ s-1
1
Expert's answer
2020-07-14T15:19:58-0400

Frequency of photon=?

Wavelength of photon("\lambda") = 1040 nm

= 1040 × 10-9 m

"\nu" = C/"\lambda"

C= velocity of light ( 2.99× 108 m/s)

Frequency= 2.99×108 m s-1 / 1040×10-9 m

Frequency of photon is

= .002875 × 1017

= 2.875 × 1014 s-1